OpenStudy (axel.caballero):

A company sells boxes of duck calls (d) for $35 and boxes of turkey calls (t) for $45. They make batches of duck calls that fill 6 boxes and batches of turkey calls that fill 8 boxes. The company only has 42 boxes. Which system of equations helps the company plan to make $300?

OpenStudy (axel.caballero):

6d = 42 - 8t 35d + 45t = 300 d = 42 - t 6d + 8t = 42 d = 42 - t 35d + 45t = 300 6d = 42 - 8t 35d - 45t = 300

OpenStudy (anonymous):

im doing the same S### I hate math